OpenStudy (anonymous):

Simplify. 4.1 + 5.6 A. 1.4 B. 1.5 C. 9.16 D. 9.7

OpenStudy (anonymous):

D

OpenStudy (anonymous):

add them together 4.1+5.6=9.7

OpenStudy (anonymous):

ok i have another

OpenStudy (anonymous):

9.7

OpenStudy (anonymous):

Simplify. –8 + 3 A. –11 B. 11 C. –5 D. 5

OpenStudy (anonymous):

9.7

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

c

OpenStudy (anonymous):

jk d

OpenStudy (anonymous):

3 more

OpenStudy (anonymous):

Simplify. 5.9 + (–3.2) A. 2.7 B. –2.7 C. –9.1 D. 9.1

OpenStudy (anonymous):

because if the first number is a negative and the second one is a positive it will always turn to a negative and a negative with a negative turns positive

OpenStudy (anonymous):

tell me the answer then you get a medal

OpenStudy (anonymous):

second one is -5 third one is 2.7

OpenStudy (anonymous):

unless the fist number is positive then negative

OpenStudy (anonymous):

2.7

OpenStudy (anonymous):

5.9 + (-3.2) = 2.7 is the same as 5.9 - 3.2 = 2.7